OpenStudy (anonymous):

For the reaction A+2B→C under a given set of conditions, the initial rate is 0.100 M/s What is change in B over the change in T under the same conditions?

OpenStudy (aaronq):

\(rate=-\dfrac{1}{1}\dfrac{[A]}{\Delta t}=-\dfrac{1}{2}\dfrac{[B]}{\Delta t}=\dfrac{1}{1}\dfrac{[C]}{\Delta t}\)